OpenStudy (anonymous):

Jake has two similar cylindrical pipes. The radius of the first cylindrical pipe is 5 cm. The circumference of the second cylindrical pipe is 20π cm. The volume of the second cylindrical pipe is how many times greater than the volume of the first cylindrical pipe? Answer 3 4 5 8

OpenStudy (btaylor):

First, solve for the radius of the second pipe.. You should get r=10 (do you get how to do this?) So draw two cylinders with a 2:1 ratio.  Now calculate the volumes of the two cylinders. The smaller one has a volume of pi(x^2)(y), and the larger one has a volume of pi(4x^2)(2y) Simplify them, and you get: Cylinder 1: V=(x^2)(y)(pi) Cylinder 2: V=8(x^2)(y)(pi) So Cylinder 2 has a volume 8 times as large as Cylinder 1.